### v1

# Experimental demonstration of spontaneous symmetry breaking with emergent multi-qubit entanglement

Abstract
Spontaneous symmetry breaking (SSB) is crucial to the occurrence of phase transitions. Once a phase transition occurs, a quantum system presents degenerate eigenstates that lack the symmetry of the Hamiltonian. After crossing the critical point, the system is essentially evolved to a quantum superposition of these eigenstates until decoherence sets in. Demonstrations of SSB with quantum-mechanical characters have been observed in superradiant phase transitions with a single qubit coupled to a single photonic mode. We here propose and demonstrate SSB in the quantum phase transition of a multi-qubit system, governed by the competition between individual drivings and intra-qubit interactions. The quantum-mechanical SSB is manifested by the emergence of a multi-qubit cat state, formed by two degenerate symmetry-breaking eigenstates. We synthesize the theoretical model in a circuit quantum electrodynamics system, where 6 Xmon qubits are coupled in an all-to-all manner through virtual photon exchange mediated by a resonator. The observed nonclassical features go beyond the conventional description of SSB, shedding new light on phase transitions for quantum many-body systems.
The idea of spontaneous symmetry breaking (SSB) lies at the heart of the Landau-Ginzburg-Wilson paradigm for describing phase transitions [1]. This idea was first proposed in solid physics, and is crucial to the understanding of phenomena in condensed matter physics, e.g., superconductivity [2]. In the 1960s, the SSB mechanism of the BCS theory of superconductivity was transported to particle physics [3, 4, 5, 6, 7], successfully explaining the origin of the mass of non-Abelian gauge bosons carrying weak forces. This breakthrough rendered the Yang-Mills gauge theory [8] to become a powerful framework for the foundation of the Standard Model [9], where the electromagnetic, weak, and strong interactions are unified. SSB also plays an important role in the evolution of the universe. During the expansion and coolingdown of the universe, it might have undergone SSB phase transitions, which transformed the “vacuum” from a state with higher symmetries to one with lower symmetries [10, 11].
The most typical example for illustrating SSB is a classical model, where a point-like particle moves along a straight line under the influence of a sombrero potential, given by $V(x)=\Lambda x^{4}-\mu x^{2}$ with $\Lambda,\mu>0$ [12]. The particle has two different positions of stable equilibrium, $x=\pm\sqrt{\mu/2\Lambda}$. Although the potential possesses a global spatial-inversion symmetry, i.e., $V(-x)=V(x)$, each of the two equilibrium positions loses this symmetry. Which equilibrium position the particle chooses to stay at depends on perturbation. This is the standard understanding of SSB, which, however, does not grasp the quantum-mechanical feature of SSB. For quantum systems, the superposition principle allows degenerate stable symmetry-breaking eigenstates to be superimposed with each other, forming a symmetric one. As such, SSB is said to occur if there exist two or more degenerate eigenstates, each of which lacks the symmetry of the governing Hamiltonian [12]. For an isolated quantum system initially with the same symmetry as the Hamiltonian, its symmetry will not be changed. To keep this symmetry, the system should evolve to the symmetric superposition of the corresponding degenerate eigenstates after a quantum phase transition. For a real quantum system, it is the environment-induced decoherence that destroys the quantum coherence, forcing the system to choose one of the symmetry-breaking eigenstates as the reality [13, 14, 15].
During the past few years, quantum phase transitions have been probed in different many-body systems, including neutral atom arrays [16, 17], trapped ions [18, 19, 20, 21], superconducting qubits [22, 23, 24], and ultracold atoms trapped in optical cavities [25, 26, 27, 28, 29, 30, 31]. In particular, symmetry-breaking phases have been experimentally explored in Dicke-like quantum phase transitions [26, 27, 28]. However, the driven-dissipative character of the realized phase transitions makes it impossible to observe the quantum coherence between the SSB degenerate eigenstates. So far, experimental demonstrations of SSB with quantum-mechanical features have been confined to a photonic mode that is coupled to a single qubit [32, 33]. For a finite many-qubit system, one of the most essential characteristics of the SSB phase transition is the appearance of a multi-qubit maximally entangled state [34, 35, 36], also referred to as the Greenberger-Horne-Zeilinger (GHZ) state [37]. Despite being fundamentally important, such a highly nonclassical SSB feature has not been experimentally observed.
We here propose and demonstrate an alternative scheme for observing SSB in the $N$-qubit Lipkin-Meshkov-Glick (LMG) model [38], where each qubit is subjected to a continuous drive and coupled to all other qubits. By slowly tuning the driving parameter, the system can evolve from an initial symmetric product state to a superposition of two degenerate symmetry-breaking eigenstates of the Hamiltonian, as a consequence of the SSB. We demonstrate this SSB process in a superconducting circuit, where 6 qubits are nearly homogeneously coupled to one another with the assistance of a bus resonator. We investigate how the symmetry is broken during the quench, and verify the quantum entanglement of the output state by measuring multi-qubit quantum correlations and mapping out the Wigner function. Our results reveal the fundamental distinction between SSB of quantum multipartite systems and that of classical ones.
The system under consideration corresponds to the $N$-qubit LMG model, featuring the competition of the individual drivings and qubit-qubit swapping couplings. The system dynamics is governed by the Hamiltonian (setting $\hbar=1$)
$H=-\Omega S_{x}-\lambda S_{z}^{2},$ (1)
where $\Omega$ is the driving strength, $\lambda$ denotes the qubit-qubit coupling, $S_{x/z}=\frac{1}{2}\stackrel{{\scriptstyle N}}{{\mathrel{\mathop{\sum}\limits_{j=1}}}}\sigma_{j}^{x/z}$ is the collective angular operator, and $\sigma_{j}^{x/z}$ denotes the Pauli operators of the $j$th qubit. The Hamiltonian has $Z_{2}$ symmetry, defined as the joint parity: ${\cal P}_{x}=\otimes_{j=1}^{N}\sigma_{j}^{x}$. The system’s ground-state behavior depends on the competition of longitudinal couplings and transverse drivings. When $\Omega,\lambda>0$ and $\Omega\gg\lambda$, the longitudinal couplings are suppressed, so that the system presents a unique ground state of the form $\left|X\right\rangle=\otimes_{j=1}^{N}\left|+_{j}\right\rangle$, where |±$\left|\pm_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle\pm\left|e_{j}\right\rangle)$. Here $\left|g_{j}\right\rangle$ and $\left|e_{j}\right\rangle$ denote the lower and upper levels of the $j$th qubit. This eigenstate has the same symmetry as the Hamiltonian, i.e., ${\cal P}_{x}\left|X\right\rangle=\left|X\right\rangle$. When $\Omega\ll\lambda$, the transverse driving only slightly shifts the energy levels of the eigenstates $\left|J,m\right\rangle$ ($\left|J,m\right\rangle$ is the eigenstate of $S_{z}$ with eigenvalue $m$ and total angular momentum quantum number $J$. See Supplementary Material [39] Sec. S1 for the detailed derivation). The most remarkable feature of the system in the strong-coupling regime is that it has two degenerate ground states, $\left|J,J\right\rangle=\otimes_{j=1}^{N}\left|e_{j}\right\rangle$ and $\left|J,-J\right\rangle=\otimes_{j=1}^{N}\left|g_{j}\right\rangle$, each breaking the system’s $Z_{2}$ symmetry.
Suppose that the system is initially in the state $\left|X\right\rangle$, which is the lowest eigenstate of the Hamiltonian under the condition $\lambda/\Omega\rightarrow 0$. When the control parameter $\Omega$ is slowly decreased, the SSB will occur. In the limit $\Omega\ll\lambda$, the system will be moved to the symmetric superposition of the two degenerate ground states, $\left|\psi_{+}\right\rangle=\frac{1}{\sqrt{2}}(\left|J,J\right\rangle+\left|J,-J\right\rangle)$, which corresponds to an $N$-qubit GHZ state. As the slow variation of the transverse drive does not change the symmetry of the Hamiltonian, the simultaneous emergence of the two symmetry-breaking eigenstates is a result of SSB quantum phase transition. The SSB is manifested by the vulnerable entanglement properties of the GHZ state, which will be transformed into a classical mixture by the environment-induced decoherence within a timescale $\tau_{\rm dec}\sim 1/N$ [15]. We note that the adiabatic preparation of superpositions of two degenerate ground states has been demonstrated in Rydberg atom arrays [40], but where the emergence of the degeneracy cannot be attributed to SSB as the Hamiltonian itself has no symmetry.
The LMG model can be engineered with a circuit quantum electrodynamics (QED) architecture involving $N$ qubits with the same frequency $\omega$, each interacting with a microwave resonator with a fixed frequency $\omega_{b}$, and subjected to a continuous drive with the frequency $\omega_{o}$, the strength $\Omega$, and the phase $\pi$. Under the condition that $\omega-\omega_{b}$ is much larger than the qubit-resonator coupling strength $\xi$, the resonator will not exchange photons with the qubits, but can induce couplings between any pair of qubits through virtual photon exchange [41, 42, 43, 44, 45]. When the resonator is initially in the vacuum state, it will remain in this state throughout the interaction, so that it can be discarded in the description of the state evolution of the qubits. Discarding a trivial constant term, the resonator-induced qubit-qubit couplings are described by the effective Hamiltonian $\lambda(-S_{z}^{2}+S_{z})$, where $\lambda=\xi^{2}/|\omega_{o}-\omega_{b}|$. When the frequency of the continuous drive satisfies $\omega_{o}=\omega+\lambda$, the system Hamiltonian in the framework rotating at the frequency $\omega_{o}$ is described by Eq. (1).
Figure 1: Experimental realization of SSB in a multi-qubit system. (a) Device schematic. The device comprises 10 Josephson-junction-based qubits and a bus resonator. The first 6 qubits, denoted as $Q_{j}$ ($j=1$ to $6$), are approximately homogeneously coupled to the resonator with the on-resonance photonic swapping rate $\xi\simeq 2\pi\times 20$ MHz. The other four qubits are unused in the experiment (not shown here). (b) Synthesis of the LMG model. When these qubits are detuned from the resonator by the same amount $|\Delta|\gg\xi$, any pair of qubits, $Q_{j}$ and $Q_{k}$, are coupled through virtual photon exchange mediated by the resonator, with the effective coupling strength $\lambda\simeq\xi^{2}/|\Delta|$. Each of these qubits is driven by a continuous microwave with the Rabi frequency $\Omega$. The system ground state depends upon the competition between resonator-induced qubit-qubit couplings and transverse drivings. (c) Pulse sequence. The experiment starts by tuning the qubits to the corresponding idle frequencies, where they are prepared in the state $\otimes_{j=1}^{6}\left|+_{j}\right\rangle$ with a continuous microwave pulse. By tuning all the qubits to the same operation frequency and applying continuous drivings, the system dynamics is described by the LMG Hamiltonian. The subsequent quench process is realized by modeling the driving amplitudes as $\Omega(\tau)=\Omega^{0}\exp(-\tau/t_{f})$, where $\Omega^{0}/(2\pi)=40$ MHz, $t_{f}=60$ ns, $\tau$ represents the quench time. Following a preset quench process, the qubits are biased back to their idle frequencies for the state readout.
Our circuit QED system is composed of 10 frequency-tunable superconducting qubits ($Q_{j}$), coupled to a bus resonator with a fixed frequency $\omega_{b}/(2\pi)\simeq 5.796$ GHz. The LMG model is realized with the first 6 qubits that are almost homogeneously coupled to the resonator with the photonic swapping rate $\xi\simeq 2\pi\times 20$ MHz, as sketched in Fig. 1(a). The other 4 unused qubits (not shown) are highly detuned from the resonator and the first 6 qubits, so that they do not affect the engineered LMG model. The energy decaying times ($T_{1}$) of these qubits range from 14 to 28 $\mu$s, while the dephasing times ($T_{2}$), measured at the corresponding idle frequencies, range from 2 to 5 $\mu$s. The system parameters are detailed in the Supplementary Material [39] Sec. S2 (see also Ref. [44]). As the qubits’ maximum frequencies in our system are close to $\omega_{b}$, the LMG model is synthesized at the operation frequency $\omega_{o}/(2\pi)=5.6895$ GHz, which is lower than $\omega_{b}/(2\pi)$ by an amount of 106.5 MHz [Fig. 1(b)]. The resulting intra-qubit couplings induced by the virtual photon exchange have a magnitude of $\lambda=2\pi\times 3.8$ MHz. The combination of these effective swapping couplings and transverse drivings with a zero phase is described by the effective Hamiltonian $-H$. The SSB quantum phase transition associated with the highest eigenstate of $-H$ is equivalent to the ground phase transition of $H$. Before the experiment, all qubits are initialized to their ground states at the idle frequencies. The pulse sequence for realizing the quantum phase transition is shown in Fig. 1(c). The experiment starts by tuning all these qubits to their idle frequencies. Then a $\pi/2$ pulse is applied to each qubit, transforming it to the superposition state $\left|+_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle+\left|e_{j}\right\rangle)$. This is followed by tuning all the qubits to the operation frequency and applying continuous drivings, incorporating the transverse drivings with the longitudinal couplings. The phase transition is realized by a quench process, during which the driving strengths are modeled as $\Omega(\tau)=\Omega^{0}\exp(-\tau/t_{f})$, where $\Omega^{0}/(2\pi)=40$ MHz, $t_{f}=60$ ns, and $\tau$ represents the quench time, as shown in Fig. 2(a). After a preset quench time, the transverse drives are switched off, following which each of the qubits is biased back to its idle frequency for the state readout.
Figure 2: Observed populations evolutions. (a) $\Omega$ versus $\tau$ during the quench process. (b) Measured population, $P_{0}^{e}$, as a function of the quench time. (c), (d), (e), (f) Population distributions measured at $\tau=0$ ns (c), $\tau=50$ ns (d), $\tau=100$ ns (e), and $\tau=150$ ns (f). For clarity, we combine all populations with the same excitation number $n$ into a single one, denoted as $P_{n}^{e}$.
To characterize the state evolution, it is necessary to be able to distinguish between $\left|\psi_{0}\right\rangle=\left|+_{j}\right\rangle^{\otimes 6}$ and $\left|\psi_{1}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle^{\otimes 6}+\left|e_{j}\right\rangle^{\otimes 6})$ that are the ground states in the limits $\lambda/\Omega\rightarrow 0$ and $\Omega/\lambda\rightarrow 0$, respectively. Though both states possess a zero z-axis magnetization $\sum_{j=1}^{N}\left\langle\sigma_{j}^{z}\right\rangle$, they have very different population distributions. For $\left|\psi_{0}\right\rangle$, the probability of all qubits being populated in $\left|g\right\rangle$ is $P_{0}^{e}=1/2^{6}$, while that for $\left|\psi_{1}\right\rangle$ is $1/2$. The measured population, $P_{0}^{e}$, as a function of the quench time, is displayed in Fig. 2(b). As expected, $P_{0}^{e}$ is increased monotonously with the quench process, during which $\Omega/\lambda$ is decreased. To reveal the system evolution more detailedly, we display the population distributions measured at $\tau=0$, $\tau=50$, $\tau=100$, and $\tau=150$ ns in Figs. 2(c), 2(d), 2(e) and 2(f), respetively, which confirm the system population undergoes a significant change during the quench, transformed from a binomial distribution to a pattern dominated by the two maximally distinct components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. Each of these components corresponds to a symmetry-breaking order with respect to the z-axis magnetization. We note that the dynamics of the qubit chain deviates from a unitary process due to dissipation.
These populations do not reflect the quantum coherence between $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. This coherence can be characterized by the nonvanishing quantum correlations among the transverse components of the Bloch vectors, denoted as
${\cal C}_{6}^{TQ}=\left\langle(\sigma_{j}^{\beta})^{\otimes 6}\right\rangle-\left\langle\sigma_{j}^{\beta}\right\rangle^{\otimes 6},$ (2)
where $\sigma_{j}^{\beta}=\cos\beta\sigma_{j}^{x}+\sin\beta\sigma_{j}^{y}$. To measure ${\cal C}_{6}^{TQ}$, before the state readout, each qubit is subjected to a $\pi/2$ rotation around the axis with an angle $\beta-\pi/2$ to the x-direction on the equatorial plane. The off-diagonal element of the system density matrix, $\rho_{e,...,e;g,...,g}$, is directly related to ${\cal C}_{6}^{TQ}$, by ${\cal C}_{6}^{TQ}=2\left|\rho_{e,...,e;g,...,g}\right|\cos(6\beta-\gamma)$, where $\gamma=\arg\rho_{e,...,e;g,...,g}$ [44, 46]. Figure 3(a) shows ${\cal C}_{6}^{TQ}$ versus $\beta$, measured for different quench times, from which the visibility of the measured ${\cal C}_{6}^{TQ}$ as a function of $\tau$ is inferred and shown in Fig. 3(b). The results show that the quantum coherence $\left|\rho_{e,...,e;g,...,g}\right|$ does not show a monotonous behavior during the quench, which can be interpreted as a consequence of the competition between the coherent evolution and the relaxation. The Hamiltonian dynamics renders the system population to converge to $\left|g_{j}\right\rangle^{\otimes N}$ and $\left|e_{j}\right\rangle^{\otimes N}$, while the relaxation deteriorates their quantum coherence. At the beginning, the populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$ are almost 00, and so is their coherence. As the amount of coherence lost during a given time is proportional to the coherence itself, at the early stage the Hamiltonian dynamics dominates the relaxation. With the increase of the populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, the relaxation plays a more and more important role. When the decoherence-induced coherence loss exceeds the gain due to the Hamiltonian dynamics, ${\cal C}_{6}^{TQ}$ begins to descend. The sources of experimental errors include decoherence, nonadiabaticity, imperfect preparation of the initial symmetric state $\left|X\right\rangle$, limitation of the qubits-resonator detuning, and the nonhomogeneous couplings (see Supplemental Material [39] Sec. S4 for details). As dephasing noises are suppressed by the transverse drivings [47], the energy relaxation is the main decoherence effect.
Figure 3: Measured quantum correlations. (a) 6-qubit transverse quantum correlation ${\cal C}_{6}^{TQ}$ versus $\beta$, measured for quench times $\tau=0$, 75, 105, and 150 ns. Here ${\cal C}_{6}^{TQ}$ is defined as $\left\langle\otimes_{j=1}^{6}\sigma_{j}^{\beta}\right\rangle-\left\langle\sigma_{j}^{\beta}\right\rangle^{\otimes 6}$, where $\sigma_{j}^{\beta}$ is the spin component along the axis with an angle $\beta$ to the x-axis on the equatorial plane of the Bloch sphere. (b) Fringe visibility of the measured ${\cal C}_{6}^{TQ}$ as a function of $\tau$. (c) Average 2-qubit longitudinal correlation ${\cal C}_{2}^{L}$ versus $\tau$.
This transverse quantum correlation can measure the extent to which the symmetry is broken only when there are no decoherence effects, as it is zero for any classical mixture of the two symmetry-breaking components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. To quantify the symmetry-breaking extent for a general multi-qubit state, we use the average two-qubit longitudinal correlation, defined as
${\cal C}_{2}^{L}=\frac{1}{15}\sum_{k>j}\sum_{j=1}^{6}\left\langle\sigma_{j}^{z}\sigma_{k}^{z}\right\rangle.$ (3)
For the initial symmetric state, this longitudinal correlation is equal to 0. With the decrease of $\Omega$, this correlation grows, and reaches the maximum $1$ when the system is entirely populated by $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, independent of their quantum coherence. Figure 3(c) displays this correlation versus the quench time $\tau$. The result illustrates the SSB process, during which the system loses the $Z_{2}$ symmetry, as a consequence of the increasing longitudinal correlations.
The quantum coherence can be further revealed by the multi-qubit Wigner function, defined as [48, 45]
${\cal W}(\theta,\varphi)={\rm Tr}\left[\rho U(\theta,\varphi)\Pi U^{\dagger}(\theta,\varphi)\right]\text{,}$ (4)
where $U(\theta,\varphi)=\exp\left[-i\theta/2\left(\sin\varphi\sigma_{j}^{x}-\cos\varphi\sigma_{j}^{y}\right)\right]$ and $\Pi=\otimes_{j=1}^{6}(1-\sqrt{3}\sigma_{j}^{z})$. To measure the Wigner function at $(\theta,\varphi)$, the joint state readout is performed after a $\theta-$rotation around the axis with an angle $\varphi-\pi/2$ to the x-direction on the equatorial plane. The Wigner function reconstructed at $\tau=0$, 50, 100, and 150 ns are displayed in Fig. 4. The Wigner function corresponds to a quasi-probability distribution of the multi-qubit system on a unit sphere. The regions where the values are negative reflect the quantum interference effects, and the negativity serves as a signature of multi-qubit cat states [45]. The results clearly verify that the two symmetry-breaking state components are superimposed with each other during the SSB process, which represents one of the most remarkable features that distinguish phase transitions of finite-size quantum many-body systems from those of infinite-size ones. Such a signature is crucial to understand the quantum-to-classical transition in many-body systems, but has not been demonstrated so far.
Figure 4: Multi-qubit Wigner distributions measured for different evolution times. The Wigner function at $(\theta,\varphi)$ is obtained by performing a $\theta$-rotation on each qubit around the axis with an angle $\varphi-\pi/2$ to the x-direction on the equatorial plane, and then reading out its state.
In conclusion, we have presented a demonstration of the SSB quantum phase transition in the LMG model. By tuning the competition between the external driving and intra-qubit interaction, we can move the system from the $Z_{2}$-symmetric phase to the symmetry-breaking phase, where the two degenerate highest eigenstates of the governing Hamiltonian, superimposed with each other, form a cat state. We demonstrate the theoretical prediction with a 6-qubit LMG model in a superconducting processor, where each of the 6 qubits is effectively coupled to all others through virtual photon exchange, mediated by a bus resonator dispersively coupled to these qubits. By gradually decreasing the strength of individual drivings, we observe the evolution of the system from a product state into an entangled GHZ state which is equivalent to SSB in modern understanding. The nonclassicality of the produced state is confirmed and characterized by the quantum coherence measurement and multi-qubit Wigner tomography. Besides fundamental interest, the realized SSB phase transition offers a dephasing-insensitive strategy for large-scale entanglement manipulation.
We thank Yi-Zhuang You at UC San Diego for the helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12274080, No. 12204105, No. T2322030, No. T2121001, and No. 92065114), the Beiiing Nova Program (GrantNo.20220484121), and the Innovation Program for Quantum Science and Technology (Grants No. 2021ZD0300200 and No. 2021ZD0301800).